A car and a jepp are 7 km apart. The car starts from rest and travels towards the jeep with an acceleration of 3.5 m/s2. The jeep starts from rest and travels towards the car with an acceleration of 2.25 m/s2. Where and where they will meet if:
a. They started at the same time?
b. b. The jeep started ahead of the car?

Answer by Alan3354(69443)

A car and a jepp are 7 km apart. The car starts from rest and travels towards the jeep with an acceleration of 3.5 m/s2. The jeep starts from rest and travels towards the car with an acceleration of 2.25 m/s2. Where and where they will meet if:
a. They started at the same time?
------
The sum of the accelerations is 3.5 + 2.25 m/sec = 5.75 m/sec
s = at^2/2
7000 m = 5.75t^2/2
t^2 = 14000/5.75
t =~ 49.34 seconds from start
-----
Use s = at^2/2 for each vehicle to find its distance from the start point.
===================
b. The jeep started ahead of the car?
For a number, you have to know how much ahead (in time units).